### v1

# Contrastive-KAN: A Semi-Supervised Intrusion Detection Framework for Cybersecurity with scarce Labeled Data

Abstract
In the era of the Fourth Industrial Revolution, cybersecurity and intrusion detection systems are vital for the secure and reliable operation of IoT and IIoT environments. A key challenge in this domain is the scarcity of labeled cyber-attack data, as most industrial systems operate under normal conditions. This data imbalance, combined with the high cost of annotation, hinders the effective training of machine learning models. Moreover, rapid detection of attacks is essential, especially in critical infrastructure, to prevent large-scale disruptions. To address these challenges, we propose a real-time intrusion detection system based on a semi-supervised contrastive learning framework using the Kolmogorov–Arnold Network (KAN). Our method leverages abundant unlabeled data to distinguish between normal and attack behaviors effectively. We validate our approach on three benchmark datasets, UNSW-NB15, BoT-IoT, and Gas Pipeline, using only 2.20%, 1.28%, and 8% of labeled samples, respectively, to simulate real-world conditions. Experimental results show that our method outperforms existing contrastive learning-based approaches. We further compare KAN with a traditional multilayer perceptron (MLP), demonstrating KAN’s superior performance in both detection accuracy and robustness under limited supervision. KAN’s ability to model complex relationships and its learnable activation functions are also explored and visualized, offering interpretability and potential for rule extraction. The method supports multi-class classification and proves effective in safety-critical environments where reliability is paramount.
Keywords Intrusion detection $\cdot$ Contrastive learning $\cdot$ Semi-supervised learning $\cdot$ Cyber-attack detection $\cdot$ Internet of things intrusion detection $\cdot$ Kolmogorov-Arnold Network $\cdot$
1 Introduction
Today with the rapid development of networking, the internet of things (IoT) and the industrial internet of things (IIoT), including technologies such as smart home applications, medical devices, automated vehicles and sensor networks [1], the adoption of networking is accelerating. Cyber-security has become one of the most critical aspects of both cyber-space and networks. The need for secure digital environments is evident in the era of fourth industrial revolution, where internet connectivity is integrated into many aspect of modern life [2], [3], [4].
In IoT and IIoT scenarios, the intrusion detection systems are designed to identify suspicious patterns in sensor, actuator and network packet data. With the advent of machine learning (ML) and deep learning (DL), the field of intrusion detection has shifted from classical rule-based methods towards intelligent data-derive approaches. ML approaches require fewer computational resources and data compared to DL methods. However, these methods often lack generalizability needed to handle unseen data. In contrast, DL methods offer superior performance but require extensive volumes of data and much larger computational resources [2]. To address these challenges, researches have proposed various solutions. Unsupervised methods aim to address the problem of costly and time-consuming data annotation by leveraging unlabeled data [5]. However, these approaches usually lack the desired performance that is needed. For data privacy, federated learning approaches are proposed to train the networks without the need to centralize data [2]. Despite all said, many existing methods still lack the interpretability required by industries for safe and reliable operation [6].
To address data augmentation, labeled data scarcity or data imbalance handling, many algorithms rely on oversampling techniques, such as synthetic minority over-sampling technique (SMOTE) [7], or its variants. Generative appearances using generative adversarial networks (GAN) have also been explored. However, there are a few problems we can point out to rule out the use of such techniques. First, these methods will have performance degradation if the imbalance ratio is large, which is common in intrusion detection datasets. Second, in many of these datasets, the data distributions of attack and normal samples are quite similar. As a result, using oversampling techniques can lead to unclear boundaries between attack and normal cases. And regarding GANs, often the training of these networks is difficult. In this work, the use of such methods for imbalance handling and data augmentation is discouraged. Instead, we use a masking strategy to randomly mask out a certain percentage of the data.
Another commonly used technique in intrusion detection is feature selection [8]. However, one significant problem with feature selection is that if an attack targets a specific feature that has been removed during feature selection, the model may fail to detect it. In this work, we avoid feature selection, but the features with no variations in the Gas Pipeline dataset are removed. We perform no feature selection other than the mentioned one.
One of the major challenges in real-time intrusion detection is the pre-processing that could delay real-time detection of cyber intrusions. Many methods use window-based sampling approaches. the problem with such techniques is that the algorithm must wait for a complete data window, before proceeding further with operations such as averaging or applying one-dimensional convolutional neural networks (1D-CNNs) for feature extraction and final classification. Furthermore the window-based algorithms, often exhibit lower performances [9]. Another preprocessing method used frequently in the literature is principle component analysis (PCA), which reduces the dimensionality of the data or transforms the input data into a set of orthogonal components based on maximum variation along the principle components. Our approach is based on individual data samples, which eliminates the need for windowing. And the pre-processing associated with this method only consists of feature encoding and standard scaling for normalization. This preserves the interpretability of the features and timely intrusion detection.
As discussed before, the limited availability of annotated data is one of the biggest challenges in ML and DL. We mentioned the advantages and disadvantages of ML and DL. To bridge the gap between ML and DL and to address the problem of annotated data, semi-supervised approaches have been proposed where the network is first pre-trained on large amounts of unlabeled data, then using only a small labeled subset, the fine-tuning is performed. One example of these semi-supervised leaning is the SimCLR framework introduced in [10], which we adopt in this work with KAN. In this study, a feature extractor based on a single layer of KAN is trained using the abundant unlabeled data. These data points are first augmented into two different views, and the feature extractor is to create same representations for these augmented views which are literally the different views of the same data point. Once the pre-training is completed, fine-tuning is performed using only a small subset of data with labels.
To address the interpretability and explainability challenges faced by DL models, we incorporate KAN into our framework. First proposed in [11], KAN is a variant of fully connected layer, where the traditional weights are replaced with learnable activation functions in the form of splines. These splines approximate univariate functions that can capture complex patterns in the data. As will be further discussed in later sections, KAN has a better performance compared to the MLP networks with the same number of parameters. KAN decomposes the complex multi-variate inputs into multiple univariate functions, and originally was used for solving ODEs and PDEs [12], but recent studies extend KAN’s applications to time series analysis [13] and classification tasks.
The contributions of the study are listed as follows:
  • An efficient real-time intrusion detection system based on semi-supervised contrastive learning, utilizing large amounts of unlabeled data.
  • Utilizing a small fraction of labeled data and achieving performance better than other state-of-the-art methods.
  • Investigating the interpretability of the proposed contrastive-KAN method through visualization of the learned spline activation functions.
  • A thorough ablation study and comparison to traditional MLP networks that explores different hyper-parameters of the KAN model to determine which hyper-parameters have the most impact on the classification results.
The remainder of the paper is organized as follows: section 2 discusses the related work in the literature, section 3 provides a detailed description of the datasets, key characteristics, the features removed, and data ratio, section 4 outlines the problem formulation, the Kolmogorov-Arnold representation theory and KAN, as well as the contrastive training and fine-tuning stages, section 5 first discusses the experimental results, followed by an extensive ablation study investigating KAN’s hyper-parameters, including spline order, grid size and width of the network, visualization of the activation functions learned by the pre-trained feature extractor, and a comparative analysis with other methods are presented as well. Finally section 6 concludes the paper and summarizes the key findings.
2 Related Works
In [14], different ensemble learning approaches such as XGBoost, random forest, support vector machine (SVM) and stacked DL models, are evaluated for cyber-attack detection. This study highlights the effectiveness of the tree-based models, such as XGBoost in comparison to the DL models. The results indicate that stacked DL model exhibit greater sensitivity to outliers compared to ensemble methods. [15] also utilizes a stacked DL model for detecting cyber-attacks in SCADA systems by combining the outputs of five individual DL models. This approach achieves higher performance compared to individual models. Moreover, this method uses random forest to determine feature importance.
In [16], a CNN-LSTM hybrid network is proposed to improve the temporal dependency in the IIoT networks. However, the authors, acknowledge that the challenges of imbalance data and privacy remain unsolved, and they proposed to adopt federated learning as the future research. A probabilistic hybrid ensemble classifier (PHEC) based on federated learning for IoT networks intrusion detection, is proposed in [17]. While this framework achieves high detection rates, it appears that for the multi-class Gas Pipeline dataset, only a subset of classes was considered, not all classes. This limits fair comparison to the other state-of-the-art methods. Furthermore, the use of order reduction operations in this work, might introduce additional computational complexity. In [18], authors introduce sustainable ensemble learning framework for intrusion detection of industrial control systems (ICS), utilizing sensitivity to different attack types and historical data reuse. In this work, models receive different dynamic importance scores, based on their effectiveness in detecting different types of attacks.
A bidirectional simple recurrent unit (BiSRU) model is introduced in [19] to address the limitations of traditional recurrent neural networks (RNNs), particularly the vanishing gradient problem, using skip connections and bidirectional structure. This BiSRU model improves training efficiency and also enhances the model’s ability to capture forward and backward temporal dependencies. Although this model demonstrate effectiveness and efficiency, its performance is highly sensitive to data balance. [20] proposes an ensemble-based detection model combining deep belief networks (DBNs) with SVMs to address cyber-security challenges, by analyzing network traffic and payload features. One advantage of this method is the one-class training of the SVM model which can be performed with only normal data. However, DBN still requires a large amount of labeled data. Additionally the scalability and training time of SVM for big datasets poses a significant challenge for adoption of this framework in real-world application. The work in [21], focuses on reducing the false positive alarm, by utilizing a graph-guided masked autoencoder (GGMAE) for anomaly detection in the process industry. This approach utilizes a topology graph to simulate process models and applies the Kullback-Leibler divergence as a loss function for input-output consistency.
In [22], contrastive learning with multi-augmentation strategy is employed to encode dynamic graph embeddings. And these embeddings are subsequently passed through transformer-based module to capture evolving patterns. Then a forecasting model predicts long-term trends for anomaly detection. [23] utilizes a feature encoder that converts input sequences into hidden representations, and a neural transformation module that generates multiple latent views. The training is performed using two contrastive losses: a contextual neural contrastive loss that aligns contextual information, and a discriminative contrastive loss that ensures diversity while preserving semantic consistency. In [24], the authors propose an unsupervised hierarchical strategy-based global-local graph contrastive learning (HS-GLGCL) framework to address the challenge of limited labeled data. This approach first constructs a process topology graph based on prior knowledge, then enhances it through graph augmentation and applies graph contrastive learning to extract meaningful representations from the unlabeled data. Finally, anomalies are identified based on reconstruction error thresholds and Kullback-Leibler divergence. [25] introduces an encoder for reinterpretation learning, a decoder for latent reconstruction, a memory network for pattern recognition and a denoiser input correction, which is pre-trained by self-supervised learning. While the approach is comprehensive, the complexity of the framework can hinder real-time intrusion detection.
[26] proposes time-series anomaly detection framework based on learnable data augmentation and contrastive learning in a self-supervised setting. The main innovation of this framework is the learnable data augmentation module that creates challenging negative samples to enhance training. [27] combines contrastive learning with graph neural networks, to learn both similarities in the latent space and dependency structures between different instance pairs in a self-supervised manner.
[28] proposes an improved deviation network with feature selection, to address the need for extensive data for anomaly detection. However, relying on feature selection poses a risk of missing critical attributes, that could potentially lead to undetected intrusions. To address both challenges of limited labeled data and privacy concerns in IIoT applications, [29] utilizes semi-supervised learning integrated with federated learning. Their architecture utilizes a autoencoder with a fully connected layer for intrusions detection. This allows the use of unlabeled data while preserving data privacy. The research in [28] introduces a semi-supervised approach, incorporating information gain (IG) and PCA for feature selection, alongside a Gaussian-based deviation loss function. But feature selection in intrusion detection system, can be cumbersome for intrusion detection, as noted before.
Apart form the learning-based methods, probabilistic and mathematical approaches have also been examined in the literature. One instance is the work proposed in [30], where a probabilistically timed dynamic model for simulating physical security attacks on critical infrastructures is presented. It models an attacker’s behavior in progressing through vulnerabilities. This method calculates mission success probability based on cumulative attack time relative to a given mission duration and uses Monte Carlo sampling to account for uncertainty, and it further produces dynamic attack trees and visual flowcharts, to better represent the decision-making process. [31] employs colored Petri nets and mixed-strategy Nash equilibrium (CPNE) to predict multi-step attack paths for oil and gas intelligent pipelines. This method models asset importance and dependency relationships, and uses this model to simulate attack-defense interactions, then calculates node failure probabilities, extracting multi-step attack paths.
3 Datasets
In this section, we discuss different industrial control system (ICS) datasets, used in this research. There are plenty of datasets publicly available regarding IoT in the literature, such as SWaT [32], WADI [33]), Gas Pipeline [34], [35], UNSW-NB15 [36] and BoT-IoT [37]. There are two major type of datasets available: simulated datasets such as UNSW-NB15, and datasets collected from physical testbeds, such as SWaT, WADI, Gas Pipeline and BoT-IoT. We validated the results of our method using UNSW-NB15, BoT-IoT and gas pipeline datasets. These datasets are mainly composed of network packets. In the following, we discuss each of the datasets in detail.
3.1 UNSW-NB15 Dataset
The UNSW-NB15 dataset created by researchers at the University of New South Wales (UNSW) represents modern network traffic scenarios and includes a wide range of low-foot print attacks. A synthetic testbed was configured using the IXIA PerfectStorm tool, that generated realistic network traffic, and a variety of attack scenarios. This dataset contains network traffic collected from multiple sources, including safe communications, malicious attacks, and routine operations such as software updates. Raw network packets were captured using the tcpdump tool and stored in PCAP file format.
A total number of 49 network-based features were extracted in five categories: basic features, content features, time features, connection features, and additional features, such as protocol type, duration, packet/byte counts, payload analysis, jitter, inter-packet arrival times, and historical behavior. This dataset includes multiple modern attack families: Fuzzers, Analysis, Backdoors, DoS, Exploits, Reconnaissance, Shellcode, Generic, and Worms, in addition to Normal traffic. In this work, we utilize only a subset of categories: Normal, Fuzzers, DoS, Exploits, Generic and Reconnaissance. The total number of records is about 2.5 million, consisting of 2,218,761 normal and 321,283 attack records [36].
The UNSW-NB15 dataset was collected to address the limitations of previous datasets (e.g. KDD99/NSL-KDD datasets), such as lack of modern low-footprint attacks and outdated normal traffic patterns. Skewness and Kurtosis tests on this dataset confirm non-linearity, and non-normal distributions and feature correlations that challenge traditional ML methods [38]. Older datasets like KDD99 and NSL-KDD suffered from redundant records. The UNSW-NB15 dataset addresses this issue [39]-[40]. The various classes and the corresponding number of instances for each condition are provided in Table 1.
It is worth noting that the original dataset contains 49 columns, including several categorical features: proto, service and state. We apply one-hot encoding to these features, this will result in a total of 204 columns after feature encoding. In this work, 558,954 samples are used for pre-training phase and 55,895 samples for fine-tuning stage, which corresponds to 22.01% and 2.20% of total number of samples, respectively. The fine-tuning set is sampled from the pre-training set.
3.2 BoT-IoT Dataset
The BoT-IoT dataset [37], publicly released in 2018 by the Network Security Lab at the University of New Brunswick, Canada, was collected using a smart home testbed that included various IoT devices like routers, cameras, and smart appliances. The setup was designed to mimic real-world IoT environments and includes multiple different types of cyber-attacks. Network traffic was recorded over several weeks using a passive monitoring tool called CICFlowMeter, which captured data at the network gateway.
This dataset contains network features, including protocol type, source/destination IP addresses, and port numbers, and each data point is labeled. It serves as a benchmark for evaluating intrusion detection systems, especially those focused on IoT security, and includes common IoT-related attacks. The categories present in this dataset: DDoS, DoS, Normal, Theft and Reconnaissance.
In practice, when preparing the data for training, some columns which are attack and subcategory are removed, however category is retained to serve as the target label. The proto column, being a categorical variable, is transformed using one-hot encoding, similar to preprocessing steps applied to the UNSW-NB15 dataset. After encoding, the dataset ends up with 20 features for training. For this dataset, we use 469,570 samples which are approximately 12.8% of the data for the pre-processing step, while for the fine-tuning phase 46,957 samples are employed which account for 1.28% of the total 3,668,522 samples. The fine-tuning subset is sampled from the pre-training subset of data, and we only use labels for the fine-tuning subset. Both numbers are significantly lower than the volume of data typically required by other algorithms. Table 2 presents the classes in BoT-IoT dataset along with the number of instances for each condition.
3.3 Gas Pipeline Dataset
The gas pipeline dataset is developed to study the intrusion detection in critical infrastructure systems that relies on SCADA technology. Its primary purpose is to simulate cyber-attacks targeting SCADA systems, serving as a standardized and common benchmark for evaluating different intrusion detection methods. This dataset is based on a small-scale gas pipeline testbed, where MODBUS traffic was recorded using a network data logger. The testbed is consists of an airtight pipeline connected to a compressor, a pressure sensor and a solenoid-controlled relief valve. The system regulates air pressure within the pipeline through a proportional-integral-derivative (PID) control scheme. In this research, we utilize the 2014 version of the dataset [35].
The dataset is composed of 27 columns, the result column, denotes the condition of the system. Therefore, the dataset actually has 26 actual features related to network traffic, process control and process measurements. Among these, seven features contain only a single constant value and do not contribute any useful information for the algorithm. These features are, rate, comm_write_fun, gain, reset, deadband, cycletime and crc_rate. After dropping these features, we are left with 19 features, that are going to be used for the training of the network. Regarding the labels, there are eight distinct conditions in the dataset, which are normal (Normal), Naïve Malicious Response Injection (NMRI), Complex Malicious Response Injection (CMRI), Malicious State Command Injection (MSCI), Malicious Parameter Command Injection (MPCI), Malicious Function Code Injection (MDCI), Denial Of Service (DoS) and Reconnaisance (Recon). These are denoted by labels, 0 to 7 respectively [35]. The classes along with their respective instance counts are shown in Table 3.
A total of 77,615 samples, that represents 80% of the entire dataset, are used for the pre-training phase. From this set, 7,761 samples (8% of the total data) are selected for fine-tuning. The pre-training sets require no labels, and labels are only used in the fine-tuning stage, making this approach suitable for scenarios with abundant unlabeled data.
Table 1: Number of data for UNSW-NB15 dataset
                  Category                   Number
                  Normal                   677,785
                  Fuzzers                   5,051
                  DoS                   1,167
                  Exploits                   5,408
                  Generic                   7,522
                  Recon                   1,759
                  Total                   698,692
Table 2: Number of data for BoT-IoT dataset
                  Category                   Number
                  DDoS                   1,926,624
                  DoS                   1,650,260
                  Normal                   477
                  Theft                   79
                  Recon                   91,082
                  Total                   3,668,522
Table 3: Number of data for Gas Pipeline dataset
                  Category                   Number
                  Normal                   61,156
                  NMRI                   2,763
                  CMRI                   15,466
                  MSCI                   782
                  MPCI                   7,637
                  MDCI                   573
                  DoS                   1,837
                  Recon                   6,805
                  Total                   97,019
4 Methodology
In this section, we present the methodology proposed in this study. In section subsection 4.1 begins with an overview of the Kolmogorov–Arnold representation theorem and KAN, subsection 4.2 outlines the pre-processing step associated with the method, and finally, in subsection 4.3, contrastive learning is thoroughly investigated.
4.1 KAN: Kolmogorov-Arnold Network
KAN, first introduced in [11], is an alternative to traditional fully-connected network. It exhibits strong capabilities in modeling non-linear functions and outperforms conventional MLP layers. KAN has demonstrated significant potential in a variety of domains, including, anomaly detection [41], human activity recognition [42], solving differential equations [43], time series analysis [44]. KAN has shown notable results, such as, advancement in accuracy, interpretability, and computational efficiency [11]. Before proceeding to the methodology proposed, we examine the current applications of KANs in the field of anomaly detection. [45] employs KAN for CPS attack detection, using a minimal and fast approach, which only utilizes standard-scaling, this method achieves high performance and low inference time.
In [46], the authors propose faster-KAN with the help of Gaussian radial basis functions (GRBF) to detect anomalies in IoT networks. This research is suitable for low-end devices and systems with limited computational resources. [41] introduces CCTAK, an unsupervised anomaly detection framework for IIoT networks, with a combination of CNN and KAN. The enhanced interpretability of this method is a key advantage. [47] uses KAN for credit card fraud detection. In this work, KAN outperforms the conventional MLP networks. In [48], the detection of botnet attacks is investigated. It is shown that the original KAN outperforms the traditional ML and DL methods, while also providing improved interpretability. [49] utilizes KAN for motor condition monitoring. KAN improves the real-time diagnosis of faults in the motors, achieving both better performance and interpretability compared to MLP networks. KAN is also adopted in healthcare domain. In [50], the authors performed the ECG-based cardiac abnormality detection, showcasing the effectiveness and generalizability of KAN for wearable medical devices.
KAN is grounded in the Kolmogorov–Arnold representation theorem, which initially served to solve differential equations. However, in this study, KAN is applied for classification tasks. The theorem guarantees that any multivariate continuous function $f(X)$, where $X=(X_{1},\dots,X_{T})$, can be decomposed into a sum of nested univariate functions. Meaning, for the smooth differentiable function $f:[0,1]^{T}\rightarrow\mathbb{R}^{d_{out}}$, there exist continuous univariate functions $\Phi_{q}$ and $\phi_{q,p}$, that satisfy the following equation:
$f(X)=\sum^{2T+1}_{q=1}\Phi_{q}\bigl{(}\sum^{T}_{p=1}\phi_{q,p}(X_{p})\bigr{)}$ (1)
where $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\Phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. In this formulation, each $\phi_{q,p}$ and $\Phi_{q}$ is a univariate function, and the nested sum of them enables the approximation of any continuous multivariate function. The Kolmogorov-Arnold representation theorem ensures $(2T+1)\times T$ uni-variate functions are enough to approximate any given $T$-variate function. Compared to Taylor or Fourier series approximations, that require an infinite number of terms to represent a multivariate function with high precision, KANs offer a significant advantage by approximating a function using a finite composition of learnable uni-variate functions. KANs leverage this theoretical foundation by replacing typical pre-defined activation functions in neural networks with learned uni-variate functions applied to each input dimension individually. Unlike in a traditional MLP layer, where activation functions are applied to the outputs of nodes, KAN places activation functions on the edges of the network, i.e., between connected nodes. Figure 1 visualizes a KAN with an input size of $T$ and an output size of $C$. From this point on, to model the approximated function $f(\cdot)$, we use the notation $KAN(\cdot)$. And $X^{i}\in R^{T}$ denotes a single sample, with a label $Y^{i}\in\{1,\dots,C\}$. In Figure 1, the output shape is of dimension $C$. To formulate this better, we can write that, $p=KAN(X^{i})$ where $p=[p_{1},p_{2},\dots,p_{C}]$, is a vector of class representations.
Figure 1: Visualization of a stack of $L$ KAN layers, with input dimension of $T$ and output dimension of $C$, denoting the number of classes [51].
The forward propagation of a KAN model with $L$ layers is expressed as a composition of non-linear transformations $\Phi_{q}$, each operating on the previous layer’s output:
$y=KAN(X)=(\Phi_{L}\circ\Phi_{L-1}\circ\dots\circ\Phi_{1})X$ (2)
each $\Phi_{l}$ consists of a set of learnable uni-variate functions $\phi$ applied element-wise. In contrast to traditional MLPs, where nonlinearity is introduced via fixed activation functions, KANs learn the activation functions during training. A traditional MLP network performs forward computation using linear weight matrices and fixed non-linear activations:
$y=MLP(X)=(W_{L}\circ\sigma\circ W_{L-1}\circ\sigma\circ\dots\circ W_{1})X$ (3)
here $W_{l}$ are weight matrices and $\sigma$ denotes a fixed activation function that introduces nonlinearity into the output of each layer. KANs learn the activation functions instead of using pre-defined ones. KANs use a hybrid activation function $\phi(X)$, which combines the SiLU and spline-based basis functions. The function is defined as follows:
$\phi(X)=\omega_{b}SiLU(X)+\omega_{s}Spline(X)$ (4)
where the SiLU activation function is defined by:
$SiLU(X)=\frac{X}{1+e^{-X}}$ (5)
The SiLU activation function is a continuous function, that guarantees the flow of gradient during the training process. This differentiable function has favorable optimization properties. The B-splines enable fine-grained modeling of nonlinearities through a learned basis function. Each spline function operates within a grid of size $G$:
$Spline(X)=\sum^{G+k}_{i=1}c_{i}B_{i}(X)$ (6)
here, $B_{i}(X)$ denotes a spline basis function, $c_{i}$ represent learned coefficient or control point, $G$ is the number of grid points, and $K$ denotes the spline order (the default is $k=3$ for cubic splines). The spline order $K$ determines the smoothness of the curve and higher grid size $G$ increases the resolution of the splines. As can be seen in Equation 6, a spline with order $K$ and grid size $G$ requires $G+K$ basis functions.
The total number of learnable parameters in a KAN layer is calculated as follows:
$Parameters=(d_{in}\times d_{out})(G+K+3)+d_{out}$ (7)
where, $d_{in}$ and $d_{out}$ are the input and output dimensions, $G$ is the grid size or the number of grid points, and $K$ is the spline order. Our implementation uses two stacked KAN layers. The first layer acts as the feature extractor (FE) and the second one performs the final classification. Unlike traditional MLPs, where parameter count only depends on dimensions of input and output, in KAN two other parameters, spline order and grid size, affect the number of parameters, which results in more parameters even for equivalent dimensional configurations.
4.2 Pre-processing
The only pre-processing performed in this study is just standardization with the mean ($\mu$) and standard deviation ($\sigma$) of the data. KAN has a unique ability in performing the classification of the intrusion data without the need for any sort of extra data pre-processing, which results in faster inference time. The standardization is performed as follows:
$X_{\text{scaled}}=\frac{X-\mu}{\sigma}$ (8)
we have tried min-max scaling as well, but standardization performed better than min-max scaling.
Figure 2: The contrastive pre-training framework. The data loading part involves, removing the unnecessary columns, encoding categorical features, handling NaN values, and sampling a batch of data for contrastive learning.
4.3 Contrastive learning
The contrastive learning framework was first introduced in [10], where it was applied in a self-supervised manner for visual representation learning. One of the key advantages of this method is that it can utilize large amount of unlabeled data created. Unlabeled data is common in various domains. This is a huge advantage, since annotating requires significant time, domain expertise, and financial resources. Contrastive learning provides a scalable solution for utilization of large amounts of unlabeled data. It pre-trains a FE, without the need for labels, and learns the underlying features present in the data. Once the pre-training is over, the parameters of the FE are frozen, and a final classifier is stacked to the network. The fine-tuning of the last layer is performed with just a fraction of the data that contains labels, without changing the parameters of the FE.
The process begins with the data $D=\{X_{scaled},Y\}$ where $X_{scaled}\in\mathbb{R}^{d_{T}}$ is standardized data, and also labels $Y$ are assumed to be unknown. Contrastive learning aims to learn similar representations for different views of the same data point that has gone through different data augmentations.
Since the similarity learning in the contrastive training is based on intra-batch distances, the first step involves sampling batches of size $M$ form the data $X_{scaled}$. Following the procedure, will result in $B(X_{scaled})$. The notation $B(\cdot)$ denotes a batch.
Now, we define the feature extractor denoted by $KAN_{\theta}(\cdot)$, with learnable parameters $\theta$, which takes an input data point $X$ and creates the representation $z$. As mentioned earlier that contrastive learning requires different views of the same data point, we propose a masking-based data augmentation strategy to randomly set features in the data point $X$ to zero. This masking function is defined as :
$x_{masked}=\text{Masking}(\cdot,p)$ (9)
where $p$ is a pre-defined masking percentage and it is set to $0.1$ in our case. The masking function randomly sets $p\%$ of the input features to zero. Following Equation 9 for an input batch twice, will result in two different batches of the same input $B(x_{scaled})$, which are, $B_{1}(X)$ and $B_{2}(X)$; note that for simplicity we remove the scaled keyword, that denoted the standard-scaled data. As mentioned before, the feature extractor $KAN_{\theta}(\cdot)$ must generate similar representations for corresponding elements of these two augmented batches. Let $B_{1}(z)$ and $B_{2}(z)$ represent the resulting batches generated from the two views, which $z$ denotes the representations created by the network, as follows:
$B_{1}(z)=KAN_{\theta}(B_{1}(X))$ (10)
$B_{2}(z)=KAN_{\theta}(B_{2}(X))$ (11)
where the network $KAN_{\theta}(\cdot)$, creates a batch of representations for each input batch. The input and output batches are of size $M$. The aim is to bring the representations created for corresponding elements in the augmented data closer to each other, while distancing the dissimilar representations. First, the two batches of data must be concatenate into a single batch with $2M$ samples, as follows:
$B_{contrastive}(z)=concat(B_{1}(z);B_{2}(z))$ (12)
The contrastive loss utilized in this research is based on cosine similarity:
$l_{i,j}=\log\frac{\exp(sim_{cos\_sim}(x^{i},x^{j})/\tau)}{\sum_{k=1}^{2M}1_{k\neq i}\exp(sim_{cos\_sim}(x^{i},x^{k})/\tau)}$ (13)
where $i,j\in\{1,\dots,2M\}$, denote the indices of the samples in batch $B_{contrastive}(z))$ and $sim_{cos\_sim}$ is the cosine similarity between two vectors as follows:
$sim_{cos\_sim}=\cos(\phi)=\frac{\mathbf{A}\cdot\mathbf{B}}{\|\mathbf{A}\|\|\mathbf{B}\|}$ (14)
where $\phi$ is the angle between the two latent representations. In Equation 13, the self-similarity ($i=k$) is ignored (with $1_{k\neq i}$). And $\tau$ is the temperature or scaling factor, that is set to $0.5$. $x^{i}=B_{contrastive}(z)(i)$ represents the $i^{th}$ element of the batch $B_{contrastive}(x)$, with similar definitions for $x^{j}$ and $x^{k}$.
$l$ is a $2M\times 2M$ matrix, where the elements $l(2k,2k-1)$ and $l(2k-1,2k)$ represent the pairwise similarity between the two batches. The total loss is then calculated as:
$Loss=\frac{1}{2M}\sum_{k=1}^{M}(l_{2k-1,2k}+l_{2k,2k-1})$ (15)
Since, the denominator of Equation 13 includes all the other dissimilar samples, it captures the dissimilarity associated with the negative samples. Maximizing the distance between the dissimilar samples, will result in minimization of the loss. The parameters $\theta$ of the feature extractor are updated according to gradient descent optimization, as follows:
$\theta_{n}=\theta_{n-1}-\eta\frac{\partial Loss}{\partial\theta}$ (16)
where $\eta$ is the pre-training learning rate, $n$ denotes the $n^{th}$ iteration of training. Then the weights are saved for the fine-tuning stage. Algorithm 1 and Figure 2, show an algorithmic and a visual view of the contrastive training.
Algorithm 1 Proposed contrastive learning framework
[…]
1:  Normalize data: $X_{\text{scaled}}\leftarrow\frac{X-\mu}{\sigma}$
2:  for $epoch=1$ to $epochs$ do
3:     Create a batch of training data, with size $M$: $B(X_{scaled})$
4:     Create different views of the data with augmentations:$B_{1}(X)\leftarrow\text{Masking}(B(X_{\text{scaled}}),p)$$B_{2}(X)\leftarrow\text{Masking}(B(X_{\text{scaled}}),p)$
5:     Feed each of the batches to the model $KAN_{\theta}(\cdot)$: $B_{1}(z)\leftarrow KAN_{\theta}(B_{1}(X))$$B_{2}(z)\leftarrow KAN_{\theta}(B_{2}(X))$
6:     Concatenate the two batches of representations vertically to create a batch of $2M$ samples:$B_{contrastive}(z)\leftarrow concat(B_{1}(z);B_{2}(z))$
7:     Calculate the $l_{i,j}$ with the contrastive loss, and $B_{contrastive}(z)$
8:     Compute similarity loss by adding the elements of the $(2k-1,2k)$ and $(2k,2k-1)$:$Loss\leftarrow\frac{1}{2M}\sum_{k=1}^{M}(l_{2k-1,2k}+l_{2k,2k-1})$
9:     Update feature extractor’s parameters: $\theta_{n}\leftarrow\theta_{n-1}-\eta\frac{\partial Loss}{\partial\theta}$
10:     Save model $KAN_{\theta}$
11:     Use $KAN_{\theta}$ to evaluate new data and use for fine-tuning
12:  end for
4.4 Fine-tuning
After pre-training the $KAN_{\theta}(\cdot)$, we stack another KAN layer, denoted as $KAN_{\lambda}(\cdot)$, with learnable parameters $\lambda$, to form the final classifier. From this point onward, we no longer use batches in the notation. One important aspect of the training process is that only the new layer parameters, denoted by $\lambda$, are updated, while the pre-trained parameters $\theta$ remain frozen, and will not be updated in the fine-tuning stage. The labeled dataset used during the fine-tuning is denoted as $D=(X_{scaled},Y)$, where $X_{scaled}$ is the standardized input feature vector and $Y^{i}\in\{0,\dots,C\}$ is the label. Unlike the pre-training phase, that the labels were unknown, in fine-tuning stage the labels are used during training. The notation for the $i^{th}$ data point is defined as:
$z^{i}=KAN_{\theta}(X_{scaled}^{i})$ (17)
$\zeta^{(i)}=KAN_{\lambda}(z^{i})$ (18)
$class^{(i)}=softmax(\zeta^{(i)})$ (19)
$error_{final}=\sum_{i=1}^{N}crossEntropy(class^{(i)},Y^{(i)})$ (20)
First we obtain the latent space representations from the feature extractor $KAN_{\theta}(\cdot)$, as defined in Equation 17. These representations are then passed through the newly stacked KAN layer $KAN_{\lambda}(\cdot)$ to produce the class representations, based on Equation 18. These class representations are then passed to the softmax function, which will form the class probabilities. Next, the class probabilities are fed into the cross-entropy loss function, in Equation 20 which will create the final classification loss, $error_{final}$. This loss will be used in gradient descent algorithm to update the model parameters according to:
$\lambda_{n+1}=\lambda_{n}-\alpha\frac{\partial error_{final}}{\partial\lambda}$ (21)
where $\alpha$ is the fine-tuning learning rate, and $\lambda$ is the final layer’s parameters. After completing the training process, the model can be saved and used for inference on new data. Algorithm 2, shows a pseudo-code of the fine-tuning phase of the proposed method.
Algorithm 2 Fine-tuning of the proposed intrusion detection framework
[…]
1:  Normalize with the $\mu$ and $\sigma$ derived from pre-training:$X_{\text{scaled}}\leftarrow\frac{X-\mu}{\sigma}$
2:  for each parameters $\in KAN_{\theta}$ do
[…]
4:  end for
5:  for $epoch=1$ to $epochs$ do
6:     Feed the data point $X^{i}_{scaled}$ $KAN_{\theta}(\cdot)$:$z^{i}\leftarrow KAN_{\theta}(X_{\text{scaled}}^{i})$
7:     Get class representations:$\zeta^{(i)}\leftarrow KAN_{\lambda}(z^{i})$
8:     Class probabilities: $class^{(i)}\leftarrow softmax(\zeta^{(i)})$
9:     Compute classification error: $error_{final}\leftarrow\sum_{i=1}^{N}crossEntropy(class^{(i)},Y^{(i)})$
10:     Update final KAN layer $KAN_{\lambda}(\cdot)$ parameters:$\lambda_{n}\leftarrow\lambda_{n-1}-\alpha\frac{\partial error_{final}}{\partial\lambda}$
11:     Save model $KAN_{\lambda}$
12:     Use $KAN_{\lambda}$ in conjunction with $KAN_{\theta}$, to evaluate new data and use for fine-tuning
13:  end for
4.5 Simulation Setup
The implementation of this project is based on efficient-kan implementation111www.github.com/Blealtan/efficient-kan. In this experiments, we utilize a learning rate of 0.001 for both pre-training and fine tuning stages across all datasets. The contrastive pre-training batch sizes are 64, 16, and 64 for the UNSW-NB15, BoT-IoT and Gas Pipeline datasets, respectively. And likewise, for fine-tuning, the batch sizes are 64, 32, and 2, respectively. The masking percentage is set to $10\%$ for UNSW-NB15 dataset and $20\%$ for BoT-IoT and Gas Pipeline dataset. The masking percentage for the BoT-IoT and Gas Pipeline dataset is higher, due to their lower feature dimensionality. With a smaller $p$, there is a chance of similar augmented views after performing the masking. The Adam optimizer is used during the pre-training phase, while AdamW is used for fine-tuning stage. The experiments are performed on Kaggle environment, with 32 GB of RAM, an NVIDIA Tesla P100 GPU with 16 GB of VRAM, and a Intel(R) Xeon(R) CPU 2.00 GHz.
5 Results
Refer to caption Refer to caption Refer to caption
(a) BoT-IoT (b) UNSW-NB15 (c) Gas Pipeline
Figure 3: t-SNE visualization of the feature extractor outputs, before pre-training.
Refer to caption Refer to caption Refer to caption
(a) BoT-IoT (b) UNSW-NB15 (c) Gas Pipeline
Figure 4: t-SNE visualization of the feature extractor outputs, after pre-training.
5.1 Performance Analysis
In this section, we analyze the results of the experiments on the three different datasets, described earlier. Figure 3 presents the t-SNE visualizations of the test data passed through the KAN feature extractor before pre-training and Figure 4 shows the corresponding results after completing the pre-training. These two figures demonstrate the effectiveness of the proposed contrastive learning framework. In the BoT-IoT dataset, the DoS and DDoS samples form more compact clusters and exhibit reduced overlap after the network has been pre-trained. In the UNSW-NB15 dataset, after the pre-training, the Exploits and Normal samples demonstrate lower degree of overlap, and we witness better separability between Normal and attack samples. Additionally, the Exploits and Fuzzers data points are better clustered in comparison to non-pretrained FE. The t-SNE visualizations for gas pipeline clearly confirm the positive effect of the contrastive pre-training on the separability of the test data. However, the gas pipeline dataset is the most challenging of all three, due to complex structure, smaller volume of data and fewer distinguishable features.
Figure 5: Visualization of the spline-based learned activation functions, aggregated with SiLU, for UNSW-NB15 dataset. The red points represent the control points that were mentioned in Equation 6. The observed upward trend in the curves results from the addition of SiLU function to the spline components, as seen in Equation 4.
5.2 Inference Time
We measured the inference time of the network to assess its feasibility for real-time deployment across all of datasets. We evaluated the proposed contrastive-KAN model on the entire test set and reported the average inference time per sample. For Gas Pipeline dataset the inference time was 0.41 ms; for the UNSW-NB15 dataset, it was 0.26 ms; and for BoT-IoT it was 0.13 ms. All evaluations are performed on CPU from the Kaggle environment, with the batch sizes specified during fine-tuning stage. It is notable that smaller batch sizes introduce some computational overhead, because smaller batch sizes require a larger number of iterations to process the entire dataset. With this said, the evaluation times clearly demonstrate that our proposed method is real-time, providing fast and timely intrusion detection.
5.3 Model Interpretability
Since the need for safe and reliable operation in the IIoT and IoT networks is present, explainable artificial intelligence (XAI) and model interpretability have become essential. For this purpose, KAN provides a very valuable aspect, that allows the visualization of the leaned activation functions. With this feature, engineers can extract rules form the trained intrusion detection systems, which can then be used in the rule-based approaches. This will further insure the reliability and safety of operation of the systems especially in safety-critical applications.
In Figure 5, the spline functions learned by the pre-trained FE of the UNSW-NB15 dataset, biased by SiLU activation function, are visualized. The red dots, represent the control points, discussed in Equation 6. The upward trend is due to the addition of the SiLU activation function. While there are many leaned activation functions, it’s clear that visualizing all of them is not possible. Figure 5 demonstrates the feasibility of the mentioned feature.
Table 4: UNSW-NB15 f1-scores for different combinations of spline orders and grid sizes, trained using the proposed framework. The network has two layers with dimensions [204 50 6].
Grid Size
1 5 20 50
Spline Order 1 82.03 84.08 89.43 93.66
3 83.14 87.45 90.17 91.24
10 84.01 86.25 87.35 89.31
30 83.12 85.14 87.29 88.27
Table 5: BoT-IoT f1-scores for different combinations of spline orders and grid sizes, trained using the proposed framework. The network has two layers with dimensions [20 100 2].
Grid Size
1 5 20 50
Spline Order 1 85.12 89.02 91.43 92.66
3 86.21 88.99 91.00 93.11
10 86.74 88.53 92.19 92.22
30 87.44 88.51 90.20 91.16
Table 6: Gas Pipeline F1-scores for different combinations of spline orders and grid sizes, trained using the proposed framework. The network has two layers with input [19 100 8].
Grid Size
1 5 20 50
Spline Order 1 86.15 86.70 86.32 86.10
3 86.18 86.66 86.33 86.02
10 86.27 85.86 86.33 86.17
30 86.18 85.97 86.61 86.53
Table 7: Ablation study on the UNSW-NB15 dataset using the specified data split. The grid size and spline order are derived from the best results of Table 4.
Model Architecture
Num. of
Params
F1 (%)
KAN (end-to-end) [204 50 6] 605,472 86.78
KAN (end-to-end) [204 70 6] 779,100 87.44
KAN (end-to-end) [204 100 6] 1,113,000 85.65
KAN (contrastive) [204 50 6] 605,472 93.66
KAN (contrastive) [204 70 6] 779,100 92.22
KAN (contrastive) [204 100 6] 1,113,000 91.91
MLP (end-to-end) [204 600 800 6] 608,606 83.59
MLP (end-to-end) [204 700 900 6] 779,806 84.92
MLP (end-to-end) [204 940 990 6] 1,130,236 85.15
MLP (contrastive) [204 600 800 6] 608,606 87.25
MLP (contrastive) [204 700 900 6] 779,806 88.87
MLP (contrastive) [204 940 990 6] 1,130,236 88.98
Table 8: Ablation study on the BoT-IoT dataset using the specified data split. The grid size and spline order are derived from the best results of Table 5.
Model Architecture
Num. of
Params
F1 (%)
KAN (end-to-end) [20 100 5] 137,500 90.01
KAN (end-to-end) [20 150 5] 206,250 89.27
KAN (end-to-end) [20 200 5] 275,000 89.51
KAN (contrastive) [20 100 5] 137,500 93.11
KAN (contrastive) [20 150 5] 206,250 92.41
KAN (contrastive) [20 200 5] 275,000 91.56
MLP (end-to-end) [20 220 600 5] 140,225 86.77
MLP (end-to-end) [20 330 600 5] 208,535 87.23
MLP (end-to-end) [20 440 600 5] 276,845 87.65
MLP (contrastive) [20 220 600 5] 140,225 89.89
MLP (contrastive) [20 330 600 5] 208,535 88.61
MLP (contrastive) [20 440 600 5] 276,845 89.41
Table 9: Ablation study on the Gas Pipeline dataset, using the specified data split. The grid size and spline order are derived from the best results of Table 6.
Model Architecture
Num. of
Params
F1 (%)
KAN (end-to-end) [19 100 8] 21,600 83.23
KAN (end-to-end) [19 150 8] 32,400 84.14
KAN (end-to-end) [19 200 8] 43,200 84.25
KAN (contrastive) [19 100 8] 21,600 86.70
KAN (contrastive) [19 150 8] 32,400 85.96
KAN (contrastive) [19 200 8] 43,200 86.57
MLP (end-to-end) [19 100 180 8] 21,628 77.57
MLP (end-to-end) [19 155 180 8] 32,628 78.92
MLP (end-to-end) [19 150 250 8] 42,758 79.12
MLP (contrastive) [19 100 180 8] 21,628 80.13
MLP (contrastive) [19 155 180 8] 32,628 81.78
MLP (contrastive) [19 150 250 8] 42,758 82.21
5.4 Ablation Study
For the ablation study, we performed experiments focusing on two key KAN hyper-parameters: spline order and grid size. While KAN has multiple hyper-parameters, spline order and grid size are the most influential. The results in Table 4 and Table 5 indicate that a network with higher grid size and smaller spline order manages to outperforms the cases with higher spline order. This is exactly what is mentioned in [52], which states that layers with higher grid sizes perform more robustly, despite having a higher Lipchitz constant. However, this trend does not hold for the Gas Pipeline dataset in Table 6.
Similarly we examined the performance of KAN against MLP, in two training paradigms: contrastive pre-training and end-to-end training. In Table 7-Table 9, the experiments conducted are listed. For each dataset we perform 12 experiments to determine the effect of the training paradigm and the width of the network. In these tables, we use two-layer KANs and three-layer MLP networks. For KANs, the spline order and grid size parameters are chosen from the best result of Table 4-Table 6. Under architecture columns, the first and last numbers represent the dimensionality of the input data and the number of classes that the classification is being performed on, and in MLP cases, the middle numbers indicate the dimensionality of the middle layer. And beside the architecture, the number of parameters of each network is present. For MLP, the focus is on an architecture that maintains a parameter count as in its KAN counterpart. All models, whether trained contrastively or end-to-end, are trained using the specified data splits. The difference is that, in the end-to-end setting, only the fine-tuning set is used. It is observed that KAN’s performance does not necessarily improve proportionate to the parameter count. In contrary, for MLP networks, the performance improves with increase in the number of parameters. One can deduce that tuning the hyper-parameters of KAN is much difficult compared to the parameters of MLP networks. Although the Gas Pipeline dataset is very challenging, the parameter count of models trained on Gas Pipeline is significantly lower than UNSW-NB15 and BoT-IoT.
It is observable that the KAN outperforms the MLP networks, that are tailored to have the same number of parameters as the KAN. In the end-to-end training, the networks are trained using the same fine-tuning set, and the pre-training with unlabeled data is not performed. The contrastive training, taking advantage of the large volumes of unlabeled data, further improves the performance.
It is worth noting that, for KANs, the width of the network has little impact on the performance. As can be observed, the networks with width of 50 or 100, outperforms cases with widths 150 and 200. This supports the earlier findings that the main factors influencing the performance of KAN are the grid size and spline order. However, for MLP networks, increasing the width of the networks, often results in marginally better performance. This indicates that, for MLP networks, there is a more predictable relationship between network capacity and the performance, or in other words, MLP networks are easier to tune.
Table 10: Comparison of our proposed methods with other approaches, UNSW-NB15 dataset.
Literature Method Accuracy (%) Precision (%) Recall (%) F1 (%)
[53] Artificial Immune System 85.00 89.00 86.00 87.00
[53] Euclidean Distance Map 90.00 92.00 91.00 91.00
[53] Filter-based SVM 92.00 94.00 92.00 93.00
[54] Decision Tree 97.08 - - 93.21
[54] Logistic Regression + RQA 95.75 - - 90.02
Proposed Method Contrastive-KAN 98.69 89.43 95.85 93.66
Table 11: Comparison of our proposed methods with other approaches, BoT-IoT dataset.
Literature Method Accuracy (%) Precision (%) Recall (%) F1 (%)
[55] Deep Neural Network 97.57 - - 93.00
[55] Support Vector Machine 83.86 - - 42.00
[55] Random Forest 92.21 - - 56.00
[56] Convolutional Neural Network 97,01 - - -
[57] Recurrent Neural Network - 98.20 85.40 88.80
Proposed Method Contrastive-KAN 97.52 90.41 94.72 93.11
Table 12: Comparison of our proposed methods with other approaches, Gas Pipeline dataset.
Literature Method Accuracy (%) Precision (%) Recall (%) F1 (%)
[58] Long-Short Term Memory 87.00 97.00 59.00 73.00
[58] Packet Signature + LSTM 92.00 94.00 78.00 85.00
[59] Bayesian network - 84.40 85.10 84.60
Proposed Method Contrastive-KAN 91.99 80.16 96.85 86.70
5.5 Comparative Analysis
Table 10, Table 11 and Table 12 show a comparative analysis of the state-of-the-art. The performance of the propose method is superior compared to other learning-based algorithms, some of which utilize more labeled data than the proposed method. This proves that the proposed algorithm is a good suit for scenarios with label scarcity.
6 Conclusion
In this paper, we proposed a semi-supervised contrastive learning approach for real-time intrusion detection under limited labeled data conditions. This framework demonstrate strong performance on IoT and IIoT datasets, outperforming other state-of-the-art methods. The results produced by this study shows that the performance of the KAN is superior to that of the MLP networks, when labeled data is scarce. Additionally, the contrastive training improves the performance of the network by utilizing the available unlabeled data. A key advantage of the proposed method is the the interpretability provided by KAN, in form of spline function visualization. This may be utilized to extract relevant rules from the network, which is beneficial in safety-critical applications, that further increases reliability, stability, and ensuring safe operation. One of the findings of this paper is that the spline order and grid size are the most influential KAN hyper-parameters to tune, while the layer dimensions have minimal impact on improving the performance. For our datasets, a higher grid size and lower spline order yields better performance. The other key finding, is that the hyper-parameter tuning for MLP networks are much easier compared to their KAN counterpart, since KAN does not exhibit a straightforward relationship between performance and the number of parameters or the increase in spline order and grid size. Although this study performed well in intrusion detection problems with limited labeled data, further studies must be performed on KAN architecture for an unsupervised case, where the training requires no labeled data, which is left for future work.